# Galvanometer-scanning transient phase microscopy with balanced detection and arbitrary pump polarization

### v2

@@ 1 Introduction
Here, we extend TM with birefringent inline interferometry to a galvanometer-scanning microscope, imaging biological and crystalline samples in a transmission geometry. With this setup we are able to compare two primary differences between TAM and TM: (1) differences in the SNR for a particular sample at our set pump and probing wavelengths and (2) additional sample information revealed by instantaneous phase interactions (i.e. XPM) versus instantaneous absorption interactions (i.e. TPA), which is particularly revealing in an imaging geometry. In contrast to previous works [32, 33], we place the pump-combining dichroic after the probe-reference birefringent splitter, allowing for the freedom to arbitrarily rotate the pump polarization. We evaluate the impact that galvanometer scanning has on probe-reference re-timing and field of view, in addition to the signal-to-noise impact of polarization effects from placing the dichroic after the birefringent probe-reference splitter [37]. Furthermore, we implement a balanced detection scheme that doubles the TM signal while canceling relative intensity noise from the laser. We demonstrate spectroscopy and imaging measurements on graphene (in which TAM yields a stronger signal) and hemoglobin and red blood cells (in which TM yields a stronger signal). Our results demonstrate the feasibility of galvanometer-scan imaging using transient phase detection and suggest that the ability to switch between TAM and TM can yield an improved SNR for a given set of pump and probe wavelengths.
@@ 2.1 Experimental Setup
The system diagram after the laser source is shown in Fig. 1. A motorized stage delays the probe relative to the pump by . A set of polarization optics (HWP 1, POL 1, HWP 2) cleans up the probe polarization and rotates it to with respect to the eigenaxes of a 2 mm thick calcite (CAL 1), with its optic axis oriented in the plane of the table. This results in an equal projection of the probe pulse along the fast (extraordinary) and slow (ordinary) axes of the crystal, respectively, resulting in two output pulses. The p-polarized reference exits first and is followed later by an s-polarized probe (see Supplementary 1A). The pump acquires a 500 kHz sinusoidal amplitude modulation by an acousto-optic modulator (AOM). A polarizer and half-wave plate provides rotatable pump polarization, before combining with the probe-reference pair in the dichroic mirror (DM) (Thorlabs DMLP900). The combined pulses are directed into a laser scanning microscope, with a pair of galvanometer scan mirrors conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm. The objective has a numerical aperture (NA) of 0.75, leading to an estimated diameter of for the pump, for the probe, and for the product of the two beams at the focus (neglecting aberrations and assuming Gaussian profiles).
@@ 2.2 Detection Technique
A diagram focusing on the polarization optics is shown in Figure 2. The probe proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1), resulting in linear p-polarized (parallel to the plane of the table) at step 1. HWP 2 rotates the polarization by (step 2). CAL 1 is a 2 mm thick piece of a-cut calcite (optic axis runs along the front face). It is oriented so that p-polarization sees the extraordinary axis and s-polarization sees the ordinary axis. Thus the probe is projected onto the ordinary ( [38]) and extraordinary axes ( [38]). Here and for are the refractive index and group index, evaluated at the center frequency of the probe pulse, . Two pulses thus exit CAL 1 (step 3) orthogonally polarized; the reference is advanced and p-polarized, the probe is delayed and s-polarized. The pair proceeds through the microscope before encountering CAL 2. CAL 2 is oriented opposite to CAL 1; the ordinary (slow) axis is rotated to coincide with the advanced reference pulse and the extraordinary (fast) axis coincides with the delayed probe pulse. The result is a re-timing of the two pulses (step 4). Note that the axes of CAL 2 may not align with s- and p-polarization as referenced to the table surface, but is instead based on the axes of the reference and probe polarizations after going through the microscope (this is due to the orientation of the detection arm which utilizes a vertical beam path relative to the tabletop). HWP 3 is used to rotate the re-timed reference and probe polarizations such that they are 45 degrees to the s- and p- axes of the polarizing beam splitter (PBS), so that both the reference and probe pulses arrive at both DET A and DET B (step 5). The pulse pair on DET B (step 6) will be out of phase from the pulse pair on DET A (step 7). Thus, if one detector measures destructive interference between the reference-probe pair then the other detector will measure constructive interference.
@@
Ultimately our setup is capable of measuring an excited-state phase change of from the initial offset point. The voltage measured on each detector is the time-integrated interference between the reference and probe; since our final signal is we end up with a 2X increase in measurable signal than if we were conducting our measurements with only one detector. The preamplifier (SR 560) measures and applies a voltage gain before the difference signal is sent into a lock-in amplifier (Liquid Instruments Moku:Lab). The lock-in amplifier measures amplitude changes of that coincide with the 500 kHz pump modulation frequency; these measurements correspond to pump-induced phase changes .
@@ 3.1.2 Graphene Monolayer
Image stacks were acquired at the edge of the graphene sample where spatially-varying structure is expected. Images of pixels were acquired with 0.05 second pixel dwell time (8.3 minutes/frame). A probe transmissivity image was acquired after setting the half-waveplate for phase measurements and the angle of incidence for the first calcite for a "balanced" state (i.e. ) with the pump blocked. From Figure 7 the edge of the sample can be seen and gives context to interpreting the image stack in Figure 8.
An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared (Fig. 8). Transient absorption still provides better SNR, but similar structural information can be seen in both.

### v3

@@ 1 Introduction
Here, we extend TM with birefringent inline interferometry to a galvanometer-scanning microscope, imaging biological and crystalline samples in a transmission geometry. With this setup we are able to compare two primary differences between TAM and TM: (1) differences in the SNR for a particular sample at our set pump and probing wavelengths and (2) additional sample information revealed by instantaneous phase interactions (i.e. cross-phase modulation, XPM) versus instantaneous absorption interactions (i.e. two-photon absorption, TPA), which is particularly revealing in an imaging geometry. In contrast to previous works [32, 33], we place the pump-combining dichroic after the probe-reference birefringent splitter, allowing for the freedom to arbitrarily rotate the pump polarization. We evaluate the impact that galvanometer scanning has on probe-reference re-timing and field of view, in addition to the signal-to-noise impact of polarization effects from placing the dichroic after the birefringent probe-reference splitter [37]. Furthermore, we implement a balanced detection scheme that doubles the TM signal while canceling relative intensity noise from the laser. We demonstrate spectroscopy and imaging measurements on graphene (in which TAM yields a stronger signal) and hemoglobin and red blood cells (in which TM yields a stronger signal). Our results demonstrate the feasibility of galvanometer-scan imaging using transient phase detection and suggest that the ability to switch between TAM and TM can yield an improved SNR for a given set of pump and probe wavelengths.
@@ 2.1 Experimental Setup
The system diagram after the laser source is shown in Fig. 1. A motorized stage delays the probe beam (prior to splitting into a reference-probe pulse-pair) relative to the pump by . (After the probe is split by CAL 1 into a reference-probe pulse-pair this temporal separation, is defined between the pump and the leading pulse from the pulse-pair, i.e. the reference pulse, inset B). A set of polarization optics (HWP 1, POL 1, HWP 2) cleans up the probe polarization and rotates it to with respect to the eigenaxes of a 2 mm thick calcite (CAL 1), with its optic axis oriented in the plane of the table. This results in an equal projection of the probe pulse along the fast (extraordinary) and slow (ordinary) axes of the crystal, respectively, resulting in two output pulses. The p-polarized reference exits first and is followed later by an s-polarized probe (see Supplementary 1A). The pump acquires a 500 kHz sinusoidal amplitude modulation by an acousto-optic modulator (AOM; TEM-100-10-1050, Brimrose Corp., Baltimore, MD). A polarizer and half-wave plate provide rotatable pump polarization, before combining with the probe-reference pair in the dichroic mirror (DM) (Thorlabs DMLP900). The combined pulses are directed into a laser scanning microscope, with a pair of galvanometer scan mirrors (QS-7, Nutfield Technology, Inc., Windham, NH) conjugated to the back focal plane of a microscope objective, followed by the sample, another objective (the condenser), and a detection arm. The objective has a numerical aperture (NA) of 0.75, leading to an estimated diameter of for the pump, for the probe, and for the product of the two beams at the focus (neglecting aberrations and assuming Gaussian profiles).
@@ 2.2 Detection Technique
A diagram focusing on the polarization optics is shown in Figure 2. The probe proceeds first through a half-wave plate (HWP 1) tuned to maximize the amount of light that proceeds through the polarizer (POL 1), resulting in linear p-polarization (parallel to the plane of the table) at step 1. HWP 2 rotates the polarization by (step 2). CAL 1 is a 2 mm thick piece of a-cut calcite (optic axis runs along the front face). It is oriented so that p-polarization sees the extraordinary axis and s-polarization sees the ordinary axis. Thus the probe is projected onto the ordinary ( [38]) and extraordinary axes ( [38]). Here and for are the refractive index and group index, evaluated at the center frequency of the probe pulse, . Two pulses thus exit CAL 1 (step 3) orthogonally polarized; the reference is advanced and p-polarized, the probe is delayed and s-polarized. The pair proceeds through the microscope before encountering CAL 2. CAL 2 is oriented opposite to CAL 1; the ordinary (slow) axis is rotated to coincide with the advanced reference pulse and the extraordinary (fast) axis coincides with the delayed probe pulse. The result is a re-timing of the two pulses (step 4). Note that the axes of CAL 2 may not align with s- and p-polarization as referenced to the table surface, but is instead based on the axes of the reference and probe polarizations after going through the microscope (this is due to the orientation of the detection arm which utilizes a vertical beam path relative to the tabletop). HWP 3 is used to rotate the re-timed reference and probe polarizations such that they are 45 degrees to the s- and p- axes of the polarizing beam splitter (PBS), so that both the reference and probe pulses arrive at both DET A and DET B (step 5). The pulse pair on DET B (step 6) will be out of phase from the pulse pair on DET A (step 7). Thus, if one detector measures destructive interference between the reference-probe pair then the other detector will measure constructive interference.
@@
Ultimately our setup is capable of measuring an excited-state phase change of from the initial offset point. The voltage measured on each detector is the time-integrated interference between the reference and probe; since our final signal is we end up with a 2X increase in measurable signal than if we were conducting our measurements with a single detector. (It should be noted, though, than both TAM and TM can be done with only one detector.) The preamplifier (SR 560) measures and applies a voltage gain before the difference signal is sent into a lock-in amplifier (Liquid Instruments Moku:Lab). The lock-in amplifier measures amplitude changes of that coincide with the 500 kHz pump modulation frequency; these measurements correspond to pump-induced phase changes .
@@ 3.1.2 Graphene Monolayer
Image stacks were acquired at the edge of the graphene sample where spatially-varying structure is expected. Images of pixels were acquired with 0.05 second pixel dwell time (8.3 minutes/frame). A probe transmissivity image was acquired after setting the half-waveplate for phase measurements and the angle of incidence for the first calcite for a "balanced" state (i.e. ) with the pump blocked. From Figure 7 the edge of the sample can be seen in this transmissivity image (left-most image) and gives context to interpreting the image stack in Figure 8. Within the transmissivity image the edge of the sample is seen at the transition from light to dark (where the light pixels correspond to the sample)
An image stack was acquired for both transient phase and transient absorption at the edge of the graphene and are compared (Fig. 8). For this sample, we see that transient absorption provides a stronger signal at overlap with both the probe and reference (-1.20 ps and 0.00 ps) than transient phase.